OpenStudy (anonymous):

35/45 points whats my grade ?

jimthompson5910 (jim_thompson5910):

(#correct/#total)*100 = grade as a percentage (35/45)*100 = 0.7556*100 = 75.56%

OpenStudy (anonymous):

thank u

jimthompson5910 (jim_thompson5910):

you're welcome